OpenStudy (anonymous):

The bill(including parts and labor) for the repair of a home appliance was $165. The cost of the parts was $85. The labor rate was 16$ per half hour. How many hours were spent in repairing the appliance?

OpenStudy (anonymous):

165=85+16x 80=16x x=(80/16) The value of X should give you how many hours

OpenStudy (anonymous):

thank you !